OpenStudy (anonymous):

if -mx^2+px-c=0 what is x equal to?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b

OpenStudy (anonymous):

c

OpenStudy (anonymous):

d

OpenStudy (anonymous):

@ranga @timo86m

OpenStudy (anonymous):

c

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

compare the given equation with the standard form ax^2+bx+c=0 here x= (-b +- sqrt(b^2 - 4*a*c)/(2a) plug a=-m b=p and c as -c

OpenStudy (anonymous):

can you help me with one more? @matricked

OpenStudy (anonymous):

may try...

OpenStudy (anonymous):

the graphs of f(x) and g(x) are shown below: what are the solutions to the equation f(x) =g(x) a. x=-3,4 b.-3.8,-0.4 c.3.8,3 d.6.5,-6

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

can you rewrite option c

OpenStudy (anonymous):

c.) -3.8, 3

OpenStudy (anonymous):

any idea on the answer?

OpenStudy (anonymous):

@matricked